the length of a rectangular platform is 2 feet longer than three times its width. the area of the platform in 56 square inches. what are the width and the length of the platform.
Found 3 solutions by ankor@dixie-net.com, ikleyn, Alan3354:
Answer by ankor@dixie-net.com(22740)

the length of a rectangular platform is 2 feet longer than three times its width.
L = 3w + 2
the area of the platform in 56 square inches.
L*W = 56
what are the width and the length of the platform
Replace L with (3w+2)
(3w+2)*w = 56
distribute w
3w^2 + 2w = 56
a quadratic equation
3w^2 + 2w - 56 = 0
You can use the quadratic formula but this will factor
(3w+14)(w-4) = 0
the positive solution is all we want here
w = 4 is the width
then
L = 3(4)+2
L = 14 is the length

Answer by ikleyn(52848)

.
the length of a rectangular platform is 2 feet longer than three times its width.
the area of the platform in 56 square inches. what are the width and the length of the platform.
~~~~~~~~~~~~~~~~~~~~


This problem is a TRAP: it has a HUGE underwater stone.

This stone is in presence of two different inconsistent units for measuring the length (feet) and the area (square inches).

So the solution by ankor@dixie-net.com is INCORRECT.



Answer by Alan3354(69443)

the length of a rectangular platform is 2 feet longer
-------------
"the length is longer"